How do I solve the following equation: (3n-6)-(5-3n)=7 ?
Answer by tiffany222(56)

(3n - 6) - (5 - 3n) = 7
3n - 6 - 5 + 3n = 7
6n - 11 = 7
6n = 18
n = 3
Final Answer: n = 3